# UU-Mamba: Uncertainty-aware U-Mamba for Cardiovascular Segmentation

### v1

@@ Abstract
Building on the success of deep learning models in cardiovascular structure segmentation, increasing attention has been focused on improving generalization and robustness, particularly in small, annotated datasets. Despite recent advancements, current approaches often face challenges such as overfitting and accuracy limitations, largely due to their reliance on large datasets and narrow optimization techniques. This paper introduces the UU-Mamba model, an extension of the U-Mamba architecture, designed to address these challenges in both cardiac and vascular segmentation. By incorporating Sharpness-Aware Minimization (SAM), the model enhances generalization by targeting flatter minima in the loss landscape. Additionally, we propose an uncertainty-aware loss function that combines region-based, distribution-based, and pixel-based components to improve segmentation accuracy by capturing both local and global features. While the UU-Mamba model has already demonstrated great performance, further testing is required to fully assess its generalization and robustness. We expand our evaluation by conducting new trials on the ImageCAS (coronary artery) and Aorta (aortic branches and zones) datasets, which present more complex segmentation challenges than the ACDC dataset (left and right ventricles) used in our previous work, showcasing the model’s adaptability and resilience. We confirm UU-Mamba’s superior performance over leading models such as TransUNet, Swin-Unet, nnUNet, and nnFormer. Moreover, we provide a more comprehensive evaluation of the model’s robustness and segmentation accuracy, as demonstrated by extensive experiments. The code can be accessed at https://github.com/tiffany9056/UU-Mamba.
keywords:
Cardiovascular Image Segmentation , Mamba , Uncertainty-Aware Loss , Sharpness-Aware Minimization
@@
[1] organization=University at Albany, State University of New York, city=Albany, postcode=NY 12222, country=USA
[…]
[2] organization=Purdue University, city=West Lafayette, postcode=IN 47907, country=USA
[…]
[3] organization=Harvard Medical School, city=Boston, postcode=MA 02115, country=USA
[…]
[4] organization=Beth Israel Deaconess Medical Center, city=Boston, postcode=MA 02215, country=USA
[…]
[5] organization=University of North Carolina at Chapel Hill, city=Chapel Hill, postcode=NC 27599, country=USA
1 Introduction
Biomedical image segmentation is crucial for medical image analysis, enabling the precise identification and delineation of anatomical structures and abnormalities [1]. Segmentation of cardiovascular structures, such as the heart, aorta, and coronary arteries, from Magnetic Resonance Imaging (MRI) and Computed Tomography (CT) is essential for diagnosing a wide range of cardiovascular conditions, developing treatment plans, and evaluating therapeutic outcomes [2, 3]. Both MRI and CT provide high-resolution images that offer detailed insights into the structure, function, and composition of these cardiovascular regions. However, manually segmenting these structures is time-consuming, labor-intensive, and prone to observer variability, underscoring the importance of automated segmentation techniques to ensure consistency and improve efficiency [4, 5].
The variability in cardiovascular anatomy, pathological changes, and the presence of imaging artifacts present significant challenges to segmenting both MRI and CT images [6]. Conventional techniques like thresholding and edge detection often fail to accurately capture the complex morphology of the heart, aorta, and coronary arteries. Recent advancements in machine learning, particularly through Convolutional Neural Networks (CNNs) [7] and other deep learning models, have shown potential in overcoming these challenges by learning intricate patterns from large datasets [8]. However, these models often require extensive computational resources and large annotated datasets, and their ability to generalize across diverse patient populations and imaging conditions may be limited [3].
To improve the generalizability and accuracy of segmentation across different cardiovascular structures, various specialized datasets have been developed. For example, the Automated Cardiac Diagnosis Challenge (ACDC) dataset [2] is designed for segmenting cardiac structures such as the left and right ventricles and myocardium from MRI. The ImageCAS dataset [9] focuses on the segmentation of coronary arteries, which is crucial for assessing coronary artery disease. The Aorta dataset [10, 11] is specifically aimed at segmenting the aorta and its branches, facilitating the diagnosis of conditions like aortic aneurysms and dissections. These datasets provide diverse challenges that contribute to the development of more robust and accurate segmentation algorithms.
Figure 1: Comparison between our method and basic approach. Traditionally, a deep learning model is trained using the cross-entropy loss $\mathcal{L}_{CE}$. Our method enhances U-Mamba by unitizing Uncertainty-aware loss $\mathcal{L}_{UA}$, which is optimized via the SAM optimizer over a flattened loss landscape. Evaluation using Dice Similarity Coefficient (DSC), Normalized Surface Dice (NSD) and Mean Squared Error (MSE) shows improvement of our method against basic CNN-based methods.
In this paper, we tackle these challenges by introducing the UU-Mamba model, an enhanced version of the U-Mamba model [12]. The UU-Mamba model incorporates a novel uncertainty-aware loss function alongside the Sharpness-Aware Minimization (SAM) optimizer [13], improving both training stability and performance. Our uncertainty-aware loss function is built upon three key components inspired by a survey of various loss functions for semantic segmentation tasks [14]:
  • Region-based loss: This component is utilized for object detection and localization, such as the dice loss [15].
  • Distribution-based loss: This component compares the predicted distributions with the ground truth, typically using cross-entropy loss [16].
  • Pixel-based loss: This component measures differences at the pixel level, employing techniques such as focal loss [17].
Instead of fixed weights, our model can dynamically modify the contribution of each loss component in accordance with the uncertainty of each prediction by employing auto-learnable weights [18]. This enables the model to prioritize predictions that are certain while simultaneously reducing the impact of ambiguous or noisy data. The auto-learnable weights allow the model to adaptively balance various aspects of the segmentation task, resulting in enhanced overall performance and robustness, particularly in mitigating challenges such as class imbalance [19]. The Sharpness-Aware Minimization (SAM) optimizer [13] is implemented to further improve the generalization capability of our model. SAM assists in the identification of parameter values that generate flat minima within the loss landscape, thereby enhancing the model’s generalizability [20, 21, 22, 23] and mitigating the risk of overfitting—a prevalent obstacle in deep learning applications for medical imaging [24]. Figure 1 presents a comparison of our method against existing approaches.
This paper extends our previous work by introducing the UU-Mamba model [25] for cardiac and vascular image segmentation. The following key aspects distinguish this work:
  1. 1.
    We extend the application of the UU-Mamba model from our previous work, which originally focused on cardiac segmentation, to now include vascular segmentation across multiple datasets. While the core model remains the same, we have enhanced its functionality to address the specific challenges of both cardiac and vascular segmentation. This extension demonstrates the model’s versatility and improved performance across a broader range of medical imaging tasks.
  2. 2.
    We present new and extensive results on the ImageCAS [9] and Aorta [10, 11] datasets, which were not included in our earlier work. These datasets differ not only in their imaging characteristics but also in the number of labels—while the ACDC dataset [2] used in our previous work involves only three labels, the Aorta dataset introduces a more complex scenario with 24 labels. This demonstrates the adaptability and robustness of our model to datasets with varying complexities.
  3. 3.
    To provide a more comprehensive evaluation compared to our previous work, where we used only the Dice Similarity Coefficient (DSC) and Mean Squared Error (MSE), we now also incorporate the Normalized Surface Dice (NSD) [26] and an analysis of the 3D loss landscape [27]. These combined evaluations—DSC, MSE, NSD, and 3D loss landscape—enhance our comparison and provide a more comprehensive understanding of the model’s ability to manage intricate medical imaging tasks.
2 Related Work
2.1 Cardiovascular Segmentation
The development of deep learning techniques, particularly Convolutional Neural Networks (CNNs) [7], has significantly advanced cardiovascular segmentation [28, 29], driven by comprehensive datasets like ACDC [2], ImageCAS [9], and Aorta [10, 11]. These datasets address a range of cardiovascular structures, including heart chambers, coronary arteries, and aortic branches, and present challenges related to anatomical variability and disease-specific characteristics. Multi-modality cardiac imaging segmentation, which utilizes imaging modalities like Positron Emission Tomography (PET), Single Photon Emission Computed Tomography (SPECT), MRI, and CT, aims to precisely segment anatomical structures and pathological regions. However, inherent challenges such as phase alignment, resolution, and image quality imbalances complicate the process. Traditional methods, including registration-based segmentation with multi-atlas approaches, and fusion-based segmentation techniques, address these issues by combining information across modalities [30, 31, 32, 33], but these methods are computationally expensive and often require large datasets [34, 35]. Hybrid approaches, combining both traditional and deep learning methods, are emerging to improve the robustness and clinical applicability of cardiovascular segmentation [36].
Recent advances in deep learning have improved segmentation accuracy by enabling complex representations of cardiovascular structures. For instance, CNNs like U-Net [37, 38] and its variants have been effective for cardiovascular segmentation tasks, particularly on the ACDC dataset [2]. However, these methods are prone to overfitting and issues such as class imbalance, which complicates their application across diverse datasets like ImageCAS [9] and Aorta [10, 11]. Zeng et al. [9] tackled coronary artery segmentation challenges in ImageCAS, utilizing multi-scale feature extraction to capture finer details, but the method still struggles with arteries exhibiting atypical morphology, requiring precise hyperparameter tuning.
The Aorta dataset [10, 11] focuses on segmenting the aorta and its branches, where varying aortic diameters and pathologies like aneurysms present additional complexities. Imran et al. [10] proposed the CIS-UNet, a hybrid model incorporating Context-Aware Shifted Window Self-Attention mechanisms to address spatial variability in aortic structures, though its performance remains sensitive to imaging quality and resolution. To further improve segmentation, some studies have integrated CNNs with attention mechanisms and multi-scale processing. For example, Hu et al. [39] adapted the Segment Anything Model to medical images, incorporating multi-scale processing and CNN heads for enhanced segmentation across various datasets.
A key development in segmentation accuracy came from Isensee et al. [37], who combined U-Net and V-Net architectures in the nnUNet framework [37]. While this method has proven effective, it heavily depends on extensive annotated datasets, limiting its applicability in scenarios with scarce labeled data—a frequent issue in cardiovascular imaging. To address this, transfer learning has been employed, as demonstrated by Chen et al. [33], who pre-trained models on the ACDC dataset and fine-tuned them on smaller, less-annotated datasets. Despite improving performance, transfer learning remains susceptible to domain shift issues, particularly when applied to datasets like ImageCAS and Aorta with differing data distributions.
Standard segmentation methods often rely on basic loss functions like Cross-Entropy loss, which struggle to effectively manage class imbalance or capture the finer details necessary for accurate segmentation. Recent research highlights the need to optimize for flatter minima in the loss landscape to enhance model generalization. Caldarola et al. [40] demonstrated that such optimization improves model robustness and generalization, especially when dealing with noisy or ambiguous data, which is essential for reliable cardiovascular segmentation across diverse clinical scenarios.
2.2 Mamba for Medical Image Segmentation
The Mamba architecture [41] represents a substantial advancement in medical image segmentation by integrating the capabilities of Vision Transformers (ViTs) [42] and Convolutional Neural Networks (CNNs) [7]. It is essential to achieve precision in medical imaging duties by integrating global contextual information and managing long sequences [43]. U-Mamba [12] expands the conventional U-Net framework [37, 38] by integrating attention mechanisms and multi-scale processing, thereby improving the accuracy and robustness of segmentation. This is achieved by building upon this foundation. This method enables the model to concentrate on pertinent details within intricate anatomical structures and efficiently process information across a range of scales, from a broad contextual understanding to precise low-level details. Furthermore, U-Mamba incorporates deep supervision, which expedites training and enhances convergence, rendering it both efficient and dependable for clinical applications that require rapid and precise image processing. The Mamba architecture’s adaptability is further underscored by specialized variants such as Weak-Mamba-UNet [44], which are able to handle intricate scenarios with improved performance and excel in scribble-based segmentation tasks. In conclusion, models that are based on U-Mamba exhibit superior segmentation performance in a variety of medical applications, such as histopathological imaging and cardiac MRI.
The Mamba architecture’s computational efficacy is one of its major advantages, as it is the result of the strategic integration of CNNs [7] and ViTs [42]. This hybrid design capitalizes on the advantages of both architectures: ViTs’ global attention mechanisms facilitate the efficient management of intricate spatial relationships and long-range dependencies, while CNNs are adept at extracting local features with minimal computational overhead. Mamba reduces the computational burden that is typically associated with pure transformer-based models, which are resource-intensive due to their quadratic complexity in relation to the duration of the input sequence, by combining these two approaches. The hierarchical design of Mamba further improves its computational efficacy. The model is capable of processing images at various scales, enabling it to capture critical features at lower resolutions and subsequently refine these details at higher resolutions. By concentrating processing capacity in the most critical areas, this multi-scale approach minimizes the total number of computations, as opposed to implementing a uniform computation across all pixels.
Furthermore, Mamba employs efficient attention mechanisms and sparse operations to minimize the number of operations necessary for each layer. This optimization is especially advantageous in the segmentation of medical images, as the computational costs associated with high-resolution images can be prohibitive rapidly. The computational efficiency of U-Mamba is also enhanced by the deep supervision strategy, which facilitates quicker convergence during the training process. This implies that the computational resources required are further reduced by the necessity of fewer training epochs to attain high performance. Overall, the Mamba architecture optimizes the strengths of both CNNs [7] and ViTs [42] while minimizing computational demands, thereby achieving a balance between precision and efficiency. Thus, it is particularly well-suited for clinical applications that prioritize both accuracy and rapidity.
2.3 Sharpness-Aware Minimization for Medical Image Segmentation
Sharpness-Aware Minimization (SAM) [13] has gained substantial attention for its capacity to enhance the generalization of deep learning models by optimizing both the training loss and the sharpness of the loss landscape. This method has been particularly advantageous in the field of medical image segmentation, where the ability to accurately detect boundaries and generalize them across a variety of datasets is essential.
SAM has exhibited significant enhancements in model robustness and accuracy within the context of medical image segmentation. For instance, Mariam et al. [45] implemented SAM in the RF-UNet model to segment retinal vessels. In addition to enhanced metrics such as accuracy, sensitivity, and specificity, their experiments on the DRIVE dataset demonstrated a substantial decrease in both training and validation losses. This emphasizes SAM’s capacity to improve generalization and mitigate overfitting in retinal segmentation tasks.
Additional research has investigated sophisticated variations of SAM for the purpose of medical segmentation. In the context of breast ultrasound image segmentation, Hassan et al. [46] assessed a variety of sharpness-based optimizers, such as SAM. Their results suggested that SAM consistently enhanced generalization across various models, surpassing other sharpness-based optimizers such as Adaptive SAM.
Random Sharpness-Aware Minimization (RSAM) by Liu et al. [47] is another innovation in SAM that incorporates randomness to enhance the efficiency and stability of SAM optimization. RSAM’s potential for improved generalization across domains suggests that it has the potential to be effective in medical imaging, despite the fact that it has not yet been explicitly applied to medical segmentation. Li et al. [48] also proposed Friendly SAM (FSAM), which further optimizes SAM for improved performance in diverse and complex data environments. This direction could be highly relevant to medical segmentation.
Collectively, these studies demonstrate that SAM [13] and its variants are essential for the advancement of medical image segmentation, particularly in tasks that necessitate high precision and robustness, such as retinal vessel extraction, breast ultrasound segmentation, and other medical imaging challenges. The integration of SAM into medical segmentation models is expected to result in even greater improvements in generalization and performance as SAM continues to develop. The notion that SAM is the most dependable sharpness-based method in medical image analysis was further substantiated by this study.
Figure 2: Overview of our proposed UU-Mamba model: Leveraging the U-Mamba architecture, we encode input images and incorporate a novel uncertainty-aware loss function. Optimization is performed using the Sharpness-Aware Minimization (SAM) optimizer [13], which operates within a flattened loss landscape. Experiments on the ACDC dataset [2], ImageCAS dataset [9], and Aorta dataset [10, 11] perform 3D heart segmentation on cardiovascular MRI and CT images, delineating each cardiovascular labels.
3 Method
Figure 1 presents a comparison of our method against basic approaches, highlighting the advancements made by our UU-Mamba model. Figure 2 illustrates our proposed UU-Mamba architecture, showcasing improvements in the training process. This model builds upon the foundational U-Mamba structure, where input images are effectively encoded. A key innovation in our approach is the integration of a novel uncertainty-aware loss function, designed to better capture and manage the inherent uncertainties in the segmentation task. To further enhance model performance, we employ the Sharpness-Aware Minimization (SAM) optimizer [13]. This optimizer is particularly well-suited for our architecture as it operates within a flattened loss landscape, which helps in achieving more robust and generalized training outcomes. These enhancements make UU-Mamba a more effective and adaptable model for both cardiac and vascular segmentation tasks. Section §$\S$ 3.1 discusses the Mamba block and the U-Mamba network, with a focus on the integration of state space models and their effectiveness in capturing long-range dependencies. In Section §$\S$ 3.2, we introduce our uncertainty-aware loss, detailing how it combines multiple loss functions to boost model performance and robustness. Finally, Section§$\S$ 3.3 covers the Sharpness Aware Minimization Optimization, emphasizing its advantages in achieving flat minima in the loss landscape, thereby enhancing generalization and mitigating overfitting.
3.1 Mamba Block and U-Mamba Network

### v2

# An Uncertainty-Aware Generalization Framework for Cardiovascular Image Segmentation

@@ Abstract
Deep learning models have achieved significant success in segmenting cardiovascular structures, but there is a growing need to improve their generalization and robustness. Current methods often face challenges such as overfitting and limited accuracy, largely due to their reliance on large annotated datasets and limited optimization techniques. This paper introduces the UU-Mamba model, an extension of the U-Mamba architecture, designed to address these challenges in both cardiac and vascular segmentation. By incorporating Sharpness-Aware Minimization (SAM), the model enhances generalization by seeking flatter minima in the loss landscape. Additionally, we propose an uncertainty-aware loss function that integrates region-based, distribution-based, and pixel-based components, improving segmentation accuracy by capturing both local and global features. We expand our evaluations on the ImageCAS (coronary artery) and Aorta (aortic branches and zones) datasets, which present more complex segmentation challenges than the ACDC dataset (left and right ventricles) used in prior work, showcasing the model’s adaptability and resilience. Our results confirm UU-Mamba’s superior performance compared to leading models such as TransUNet, Swin-Unet, nnUNet, and nnFormer. We also provide a more in-depth assessment of the model’s robustness and segmentation accuracy through extensive experiments.
I Introduction
Biomedical image segmentation is crucial for medical image analysis, as it enables the precise identification and delineation of anatomical structures and abnormalities [zengxin2024]. Accurate segmentation of cardiovascular structures, such as the heart, aorta, and coronary arteries, from Magnetic Resonance Imaging (MRI) and Computed Tomography (CT) images is essential for diagnosing a wide range of cardiovascular conditions, developing treatment plans, and evaluating therapeutic outcomes [bernard2018deep, litjens2017survey]. Although MRI and CT provide high-resolution images that reveal detailed information about the structure, function, and composition of these cardiovascular regions, manual segmentation of these structures is time-consuming, labor-intensive, and subject to variability among observers. This limitation highlights the need for automated segmentation techniques that can improve consistency and efficiency in the analysis [petitjean2011review, maier2019gentle].
The variability in cardiovascular anatomy, pathological changes, and imaging artifacts poses significant challenges for accurate MRI and CT image segmentation [li2018]. Conventional techniques, such as thresholding and edge detection, often fail to accurately capture the complex morphology of the heart, aorta, and coronary arteries. Recent advances in machine learning, particularly through Convolutional Neural Networks (CNNs) [lecun1995] and other deep learning models, have shown promise in addressing these challenges by learning intricate patterns from large datasets [fahmy2019]. However, despite these advances, current deep learning models still require substantial computational resources and large annotated datasets, while their ability to generalize across diverse patient populations and imaging conditions remains limited [litjens2017survey].
Figure 1: Comparison between our method and basic approach. Traditionally, a deep learning model is trained using the Cross-Entropy loss $\mathcal{L}_{CE}$. Our method enhances U-Mamba by utilizing the uncertainty-aware loss $\mathcal{L}_{UA}$, which is optimized via the SAM optimizer over a flattened loss landscape. Evaluation using Dice Similarity Coefficient (DSC), Normalized Surface Dice (NSD) and Mean Squared Error (MSE) shows improvement of our method against basic CNN-based methods.
To improve segmentation accuracy and generalizability across various cardiovascular structures, several specialized datasets have been established. For example, the Automated Cardiac Diagnosis Challenge (ACDC) dataset [bernard2018deep] focuses on cardiac segmentation, including the left and right ventricles and the myocardium in MRI images. The ImageCAS dataset [zeng2023imagecas] targets coronary artery segmentation, which is essential for the evaluation of coronary artery disease. The Aorta dataset [imran2024cis, krebs2024volumetric] specifically targets the segmentation of the aorta and its branches, supporting the diagnosis of conditions such as aortic aneurysms and dissections. These datasets collectively offer diverse challenges that contribute to the development of more robust and accurate segmentation algorithms.
In this paper, we tackle these challenges by introducing an Uncertainty-Aware Generalization Framework, UU-Mamba, an enhanced version of the U-Mamba model [ma2024umamba]. The UU-Mamba model features a novel uncertainty-aware loss function along with the Sharpness-Aware Minimization (SAM) optimizer [foret2021sharpness], which improves both training stability and performance. Our uncertainty-aware loss function is composed of three key components, inspired by a review of various loss functions for semantic segmentation tasks [jadon2020]. For example, Region-based loss: Improves object detection and localization, such as Dice loss [sudre2017]. Distribution-based loss: Compares predicted distributions with ground truth, typically using Cross-Entropy loss [ma2004]. Pixel-based loss: Measures differences at the pixel level, using techniques such as focal loss [lin2017]. Instead of using fixed weights, our model dynamically adjusts the contribution of each loss component in accordance with the uncertainty of each prediction by employing auto-learnable weights [cipolla2018]. This enables the model to prioritize confident predictions while minimizing the influence of ambiguous or noisy data. The auto-learnable weights allow adaptive balancing of different aspects of the segmentation task, leading to improved overall performance and robustness, particularly in addressing challenges such as class imbalance [azad2023]. To further enhance our model’s generalization capability, we implement the Sharpness-Aware Minimization (SAM) optimizer [foret2021sharpness]. SAM helps identify parameter values that lead to flatter minima in the loss landscape, improving the model’s generalizability [lin2024robust, lin2024robust2, lin2024robust3, lin2024preserving] and reducing the risk of overfitting, a common challenge in deep learning applications for medical imaging [Mariam2024]. Figure 1 presents a comparison between our method and existing approaches.
This paper extends prior work on UU-Mamba [tsai2024uumamba] by comprehensively evaluating the model for both cardiac and vascular image segmentation tasks. The main contributions are summarized as follows:
  1. 1.
    We expand the use of UU-Mamba model, originally developed for cardiac segmentation, to include vascular segmentation across multiple datasets. While the core architecture remains unchanged, we have enhanced its capabilities to address the unique challenges of both cardiac and vascular segmentation, demonstrating the model’s versatility and improved performance across a broader range of medical imaging tasks.
  2. 2.
    New results are reported on the ImageCAS [zeng2023imagecas] and Aorta [imran2024cis, krebs2024volumetric] datasets, which were not included in prior work. These datasets differ in imaging characteristics and label complexity. For example, the Aorta dataset includes 24 labels, whereas the ACDC dataset [bernard2018deep] used previously includes three labels, enabling assessment under substantially increased segmentation complexity.
  3. 3.
    In addition to Dice Similarity Coefficient (DSC) and Mean Squared Error (MSE) used previously [tsai2024uumamba], the evaluation incorporates Normalized Surface Dice (NSD) [maier-hein2022metrics] and analyzes the 3D loss landscape [Hao2018]. Together, these provide a more comprehensive characterization of model performance and optimization behavior on complex medical imaging tasks.
II Related Work
II-A Cardiovascular Image Segmentation
The development of deep learning techniques, particularly Convolutional Neural Networks (CNNs) [lecun1995], has significantly advanced cardiovascular segmentation [wang2024artificial, zhou2019], driven by comprehensive datasets such as ACDC [bernard2018deep], ImageCAS [zeng2023imagecas], and Aorta [imran2024cis, krebs2024volumetric]. These datasets address a range of cardiovascular structures, including heart chambers, coronary arteries, and aortic branches, and present challenges related to anatomical variability and disease-specific characteristics. Multi-modality cardiac imaging segmentation, which utilizes imaging modalities like Positron Emission Tomography (PET), Single Photon Emission Computed Tomography (SPECT), MRI, and CT, aims to precisely segment anatomical structures and pathological regions. However, inherent challenges such as phase alignment, resolution, and image quality imbalances complicate the process. Traditional methods, including registration-based segmentation with multi-atlas approaches, and fusion-based segmentation techniques, address these issues by combining information across modalities [Zhuang_2019, zhao2022, eugenio2015, chen2020deep], but these methods are computationally expensive and often require large datasets [xu2023deep, Chartsias2021]. Hybrid approaches, which combine traditional and deep learning methods, are emerging to improve the robustness and clinical applicability of cardiovascular segmentation [Fu_2020].
Recent advances in deep learning have improved the precision of segmentation by enabling complex representations of cardiovascular structures. For example, CNNs such as U-Net [Isensee2021nnUNet, ronneberger2015u] and its variants have been effective for cardiovascular segmentation tasks, particularly on the ACDC dataset [bernard2018deep]. However, these methods are prone to overfitting and issues such as class imbalance, which complicates their application across diverse datasets like ImageCAS [zeng2023imagecas] and Aorta [imran2024cis, krebs2024volumetric]. Zeng et al. [zeng2023imagecas] tackled coronary artery segmentation challenges in ImageCAS, using multi-scale feature extraction to capture finer details, but the method still struggles with arteries exhibiting atypical morphology, requiring precise hyperparameter tuning.
A key development in segmentation accuracy came from Isensee et al. [Isensee2021nnUNet], who combined U-Net and V-Net architectures in the nnUNet framework [Isensee2021nnUNet]. Although this method has proven effective, it is highly dependent on extensive annotated datasets, limiting its applicability in scenarios with scarce labeled data—a frequent issue in cardiovascular imaging. To address this, transfer learning has been employed, as demonstrated by Chen et al. [chen2020deep], who pre-trained models on the ACDC dataset and fine-tuned them on smaller, less-annotated datasets. Despite improving performance, transfer learning remains susceptible to domain shift issues, particularly when applied to datasets like ImageCAS and Aorta with differing data distributions.
Standard segmentation methods often rely on basic loss functions like Cross-Entropy loss, which struggle to effectively manage class imbalance or capture the finer details necessary for accurate segmentation. Recent research highlights the need to optimize for flatter minima in the loss landscape to enhance model generalization. Caldarola et al. [caldarola2022] demonstrated that such optimization improves the robustness and generalization of the model, especially when dealing with noisy or ambiguous data, which is essential for reliable cardiovascular segmentation in diverse clinical scenarios.
Figure 2: Overview of our proposed framework, we encode input images and incorporate a novel uncertainty-aware loss function. Optimization is performed using the Sharpness-Aware Minimization (SAM) optimizer [foret2021sharpness], which operates within a flattened loss landscape. Experiments on the ACDC dataset [bernard2018deep], ImageCAS dataset [zeng2023imagecas], and Aorta dataset [imran2024cis, krebs2024volumetric] perform 3D heart segmentation on cardiovascular MRI and CT images, delineating each cardiovascular labels.
II-B Uncertainty-Aware Learning in Medical Image
Understanding uncertainty is crucial for the medical image segmentation field, where the robustness and trustworthiness of predictions are essential in practical applications. The source of uncertainty can be mainly categorized into data uncertainty and model uncertainty. While data uncertainty is irreducible, model uncertainty can be reduced by improving the data  [gawlikowski2023survey, cipolla2018]. Uncertainty-aware learning allows the model to adjust its training process based on the level of uncertainty, improving accuracy and robustness by emphasizing confident predictions and reducing the impact of uncertain ones  [cipolla2018]. Various techniques have been developed to incorporate uncertainty-aware learning. Hu et al.  [hu2023umednerf] incorporate task-dependent uncertainty into MedNeRF to reconstruct 3D CT projections from a single X-ray view. By introducing uncertainty awareness, the loss terms of different tasks are adaptively weighted based on the levels of uncertainty of each task. Their experiment on chest and knee datasets shows that uncertainty-aware automatic weighting enhances both PSNR and SSIM compared to fixed-weight baselines, demonstrating that leveraging uncertainty-aware learning enables more stable convergence and improved reconstruction quality. Xia et al.  [xia2020uncertainty] develop an uncertainty-guided co-training (UMCT) framework that utilizes Bayesian dropout to estimate model uncertainty, which is then used to weight and fuse pseudo-labels, so that predictions with higher confidence contribute more to the training. This uncertainty-aware fusion improves the reliability of supervision on unlabeled data. Nahiduzzaman et al.  [nahiduzzaman2025uncertainty] extend the Variational Information Pursuit (V-IP) framework by incorporating sample-specific uncertainty into concept-based reasoning. EUAV-IP skips unreliable concept predictions, while IUAV-IP prioritizes more reliable concepts in the query selection process, enabling the model to base its decisions on a smaller, more trustworthy subset of concepts while maintaining high diagnostic accuracy. Their models achieve improved predictive performance and shorter and meaningful explanations. Taken together, these studies indicate that uncertainty-aware learning is widely applied in medical image segmentation models and consistently outperforms conventional deterministic models under challenging conditions such as multi-task learning, limited labeled data availability, and noisy or ambiguous feature representations, by enabling models to adaptively focus on reliable information and reduce the impact of uncertain predictions.
II-C Sharpness-Aware Minimization in Medical Image Segmentation
Recent studies have demonstrated that the geometry of the loss landscape plays a critical role in determining the generalization behavior of deep neural networks. Conventional optimization methods such as stochastic gradient descent (SGD) and Adam often converge to sharp minima that fit training data well but are highly sensitive to perturbations, resulting in limited robustness on unseen data. To address this issue, Sharpness-Aware Minimization (SAM) was introduced by Foret et al. [foret2021sharpness], which explicitly seeks parameter configurations that minimize the worst-case loss within a local neighborhood of the solution. By encouraging convergence to flatter regions of the loss landscape, SAM improves model robustness and generalization. Recent theoretical analyses further confirm that SAM effectively reduces local sharpness and promotes flatter minima [wen2022does].
SAM has attracted increasing attention in medical image analysis, where robustness and generalization are particularly important due to limited annotations, heterogeneous imaging protocols, and noisy labels. In retinal vessel segmentation, Mariam et al. [iqra2024] incorporated SAM into the RF-UNet framework and observed consistent improvements in accuracy, sensitivity, and specificity, alongside reduced training–validation performance gaps. In breast ultrasound segmentation, Hassan et al. [hassan2024] conducted a systematic evaluation of sharpness-based optimizers and reported that SAM consistently improved generalization across multiple architectures, outperforming several alternative sharpness-aware variants.
Beyond standard SAM, several extensions have been proposed to further enhance optimization stability and efficiency. Random Sharpness-Aware Minimization (RSAM) introduces stochastic perturbations to reduce computational overhead while maintaining robustness [yong2022rsam], and Friendly SAM (FSAM) aims to improve optimization behavior in complex data regimes [tao2024friendly]. Although these variants have shown promising results in general vision tasks, their adoption in medical image segmentation remains limited.
Despite these advances, the role of sharpness-aware optimization has not been systematically investigated in cardiovascular image segmentation, which presents unique challenges including strong anatomical variability, thin and elongated structures (e.g., coronary arteries), complex branching patterns, and heterogeneous imaging modalities such as MRI and CT. These characteristics tend to produce highly irregular loss landscapes, making models particularly vulnerable to sharp minima and performance degradation under domain shift [caldarola2022]. This gap motivates further exploration of sharpness-aware optimization strategies tailored to cardiovascular segmentation scenarios.
III Method
III-A Overview
@@
III-B Network Architecture